OpenStudy (anonymous):

10 years, Longin Company will have to repay a $60,000 loan. Assume a 4% interest rate compounded quarterly. How much must Longin pay each period to have the $60,000 at the end of 10 years? Verify your result.

OpenStudy (tkhunny):

Do you have an accumulation formula or shall we build one?

OpenStudy (tkhunny):

\[i_{annual} = 0.04\] \[i{quarterly} = 0.04/4 = 0.01\] \[r = 1+i_{quarterly} = 1.01\] Are we getting anywhere?